Solve for the 67th term of the fibonacci sequence?


1
Expert's answer
2021-09-27T14:13:20-0400

Use Binet formula


"F_n=\dfrac{\alpha^n-\beta^n}{\alpha-\beta}"

"\alpha=\dfrac{1+\sqrt{5}}{2}, \beta=\dfrac{1-\sqrt{5}}{2}"

"\alpha-\beta=\sqrt{5}"

"F_{67}=\dfrac{(1+\sqrt{5})^{67}-(1-\sqrt{5})^{67}}{2^{67}\sqrt{5}}"

"F_{67}=27777890035288"